OpenStudy (anonymous):

Show that the line integral is independent of path, and evaluate the integral. ∫(c) y dx + (x+z) dy + y dz; C is parametrized by r(t)= (t^2+1)/(t^2-1) i +cos(pi)t j + 2tsin(pi)t k for 0<_ t <_ 1/2

OpenStudy (anonymous):

it is not related to C so just initial and final point (-1 , 1 , 0) to ((-5/3 , 0 , 2) int=xy+yz

OpenStudy (anonymous):

So thE answer is going to be 0-(-1)=1